OpenStudy (babeberry):

Solve 64x = 16x−1. x = −2 x = −1 x = negative 1 over 4 x = negative 1 over 3 I think It's D.

OpenStudy (anonymous):

ur right want a fan

OpenStudy (yuliacosta111):

are you solving for x ?

OpenStudy (babeberry):

yes

OpenStudy (math_man21):

nice bio berry

OpenStudy (babeberry):

smh thx

OpenStudy (yuliacosta111):

Ok to solve for a variable in the exponent, you want to change both sides to having the same base. So you want a common factor between 64 and 16. Which would be? :)

OpenStudy (babeberry):

I'm thinking D.

OpenStudy (math_man21):

btw i read EVERYONE'S bio

OpenStudy (math_man21):

i'd say d

OpenStudy (math_man21):

it seems to add up more than the rest

OpenStudy (babeberry):

Wait C.

OpenStudy (yuliacosta111):

4^3 = 64. I was meaning the exponent; not division :) SO it would look like this 4^3x = 4^(2x-2) Then you can set 3x=2x-2 and solve for x=-2 :) does that make sense?

OpenStudy (anonymous):

yea I think it is c too 🍕🍕🍕🍉🍉💎💎

OpenStudy (math_man21):

ohh i see it thx yuli

OpenStudy (yuliacosta111):

I think the first one

OpenStudy (yuliacosta111):

@welshfella is that right

rebeccaxhawaii (rebeccaxhawaii):